# Loops, Holonomy and Signature

### v3

@@ Abstract.
We show that there is a topology on certain groups of loops in euclidean space such that these groups are embedded in a Fréchet-Lie group which is the structural group of a principal bundle with connection whose holonomy coincides with the Chen signature map. We also give an alternative geometric new proof of Chen signature Theorem and a generalization of this theorem in classes strictly containing the one originally considered by Chen.
@@ 1. Introduction
In this paper we consider the relationship between holonomy maps of principal bundles with a connection and the signature map defined by Chen ([Ch1], [Ch2], [Ch3]). Holonomy maps appear in topology and differential geometry and they trace back to the well known theory of representations of the fundamental group and constructions of local systems over a manifold by the holonomy map of a principal bundle with a flat connection. A natural question is how to extend this construction to arbitrary connections. This question leads directly to the loop space of the manifold, an object that will be defined shortly  [^1] Do not confuse this space with the one defined in algebraic topology. The ambiguity in the names will disappear after we introduce the right notation.. In contrast to the fundamental group of a manifold, the loop space we are referring to here will not be a purely topological object. Specifically, while the fundamental group of a manifold coincides with that computed in the category of smooth curves, the loop space of the manifold, and therefore the results concerning this space, will depend on the category of curves where it is defined.
Consider a point $p$ in a manifold $M$. We say that a class of loops based at $p$ is concatenable if the concatenation of two loops in the class belongs to the class. Examples of classes with this property are the class of piecewise smooth loops $\Omega^{ps}(M,p)$, the class of piecewise smoothly immersed loops $\Omega^{psi}(M,p)$ or the class of piecewise analytic loops $\Omega^{pa}(M,p)$. Examples not having this property are the smooth and analytic classes.
@@
By identifying loops of a concatenable class by reparameterization, we obtain associativity of the operation but in general $\alpha\,\alpha^{-1}\neq c$ where $c$ denotes the constant loop. A further equivalence relation is needed. A natural relation is the one finitely generated by $\alpha aa^{-1}\beta\sim\alpha\beta$. This will be called the retrace relation. The corresponding groups are denoted by $\mathcal{L}^{\ ps,ret}(M,p)$, $\mathcal{L}^{\ psi,ret}(M,p)$ or $\mathcal{L}^{\ pa,ret}(M,p)$ depending on the choice of the class of curves.
@@
In [Tl], Tlas proves a similar result in the concatenable class of $C^{1}$ loops based at $p$ with zero derivative at $p$. He proves that if the Lie group $G$ is semisimple then a pair of loops in this class are $G$-holonomy related if and only if they are related by a rank one homotopy. In the class of piecewise smooth loops, Tlas result is equivalent to the following isomorphism under thin homotopy relation (see Remark $4$ in [Me])
Along the lines of this result, Meneses in [Me] comments that the analog result of (1) claimed in [Sp] is not valid in the class of piecewise smooth loops and he gives a counterexample of a thin and non-retreacable loop, hence
One of the difficulties with piecewise smooth loops, in contrast with piecewise analytic ones, is that they can intersect in very complicated ways. This led Baez and Sawin in [BS] to consider instead the class of piecewise smoothly immersed loops $\Omega^{psi}(M,p)$. In this class they develop the technology of tassels and webs. A careful inspection into the work in [Sp], which uses the technology developed in [BS], shows that the claimed result holds for the class of piecewise smoothly immersed loops. That is, for a connected and non-solvable Lie group $G$, the following isomorphism is proved in [Sp]  [^2] We do not know why Spallanzani did not write the piecewise immersion hypothesis.
(4) $\mathcal{L}^{\ psi,ret}(M,p)\ \cong\ \mathcal{L}^{\ psi,G}(M,p),\qquad G\ \rm{connected\ and\ non-solvable.}$
Independently, in the fifties Chen defined a map in the space of paths in the euclidean space ${\mathbb{R}}^{n}$, the signature map
(5) $\Theta(\gamma)\,=\,1\,+\,\sum^{\infty}_{p=1}\,\ \sum_{i_{1},\,i_{2},\,\ldots,\,i_{p}\,=1}^{n}\ \,\int_{\gamma}dx_{i_{1}}dx_{i_{2}}\ldots dx_{i_{p}}\ X_{i_{1}}X_{i_{2}}\ldots X_{i_{p}}$
taking values in the ring of the formal power series with the non-commutative variables $X_{1},\ldots,X_{n}$. Chen proved that this map is a faithful representation of the group of loops $\mathcal{L}^{\ psi,ret}({\mathbb{R}}^{n},p)$ [Ch1, Ch2, Ch3]. For other results concerning iterated integrals, see [Ch4, Ch5]. In recent years, this theory had a resurgence in the context of stochastic processes [Ly], bounded variation paths [HL] and rough paths [Ly, BGLY].
In this paper, we construct a principal bundle with connection whose holonomy coincides with the Chen signature map on the group of loops $\mathcal{L}^{\ psi,ret}({\mathbb{R}}^{n},p)$ in euclidean space. In other words, we show that the Chen signature map on this loop space is the holonomy map of a principal bundle with connection. In particular, this holonomy map is a monomorphism.
In view of the mentioned counterexample given by Meneses in [Me] of a non-retreacable thin loop in the class of piecewise smoothly loops showing in particular that the group $\mathcal{L}^{\ ps,ret}(M,p)$ is a proper extension of $\mathcal{L}^{\ ps,thin}(M,p)$, we show that no such counterexample is possible in the class of piecewise smoothly immersed loops. Actually, in section 3 we prove the following proposition.
@@
From now on, unless otherwise specified, we will denote the groups in the previous proposition simply by $\mathcal{L}(M,p)$ and by $\mathcal{L}_{p}$ in the case the manifold is the euclidean space ${\mathbb{R}}^{n}$. The following is the main result of the paper.
Let $\overline{G}_{n}$ be the Fréchet-Lie group corresponding to the completion of the real free Lie algebra generated by $X_{1},\ldots,X_{n}$. (See section 4 for more details and references).
@@
There is a $\overline{G}_{n}$-principal bundle with connection $(E,\,{\mathbb{R}}^{n},\,\overline{G}_{n},\,\pi,\,\nabla)$ such that for any $p$ in ${\mathbb{R}}^{n}$, the holonomy map is a monomorphism
$Hol_{\,\nabla,\,p}:\,\mathcal{L}_{p}\,\hookrightarrow\,\overline{G}_{n}$
and it coincides with the Chen signature map (5) on $\mathcal{L}_{p}$.
In [Ner], the author introduces for $\xi>0$ a family of Banach-Lie groups $\mbox{Fr}_{n}^{\,\xi}\subset\overline{G}_{n}$. The above result can be strengthened:
@@
For every loop $\gamma$ in $\mathcal{L}_{p}$, the element $Hol_{\,\nabla,\,p}(\gamma)\,=\,\Theta(\gamma)$ belongs to the Banach-Lie group $\mbox{Fr}_{n}^{\,\xi}$ for every $\xi>0$.
@@
In section 5 we give a geometric alternative new proof of the Chen signature theorem, quoted as Theorem 2.2 here, as a corollary of either the Tlas result in [Tl] or the holonomy theorem in [Sp]. A verbatim argument gives a generalization of the Chen signature theorem in the class of piecewise smooth loops under the thin homotopy relation, a strictly larger class than the one originally considered by Chen, which in turn implies the following theorem.
Theorem 1.5.
As an immediate corollary of the previous theorem, we have the following complement to the main Theorem 1.2.
@@ 2.1. Loops and Holonomy
Now, consider the class $\Omega^{psi}(M,p)$ of piecewise smoothly immersed closed curves based at the point $p$. Another equivalence relation in this class is given by the finite composition of thin homotopies. A thin homotopy between two curves $\alpha$ and $\gamma$ is a homotopy $\eta:[0,1]^{2}\rightarrow M$ such that its image is contained in the union of the images of the curves, that is
@@
Denote the quotient under this relation by $\mathcal{L}^{\ psi,thin}(M,p)$ and note that, as before, it is a group under concatenation whose neutral element is the equivalence class of $e_{p}$, the constant path at $p$.
@@
In view of this isomorphism, we will simply denote these groups by $\mathcal{L}(M,p)$.
Another possible identification is, given a Lie group $G$, identify two closed curves $\alpha$ and $\beta$ in $\Omega^{psi}(M,p)$ if they have the same holonomy for any $G$-principal bundle with connection over $M$. This relation is the $G$-holonomy relation. The group obtained this way is called the group of hoops based at $p$ and will be denoted by $\mathcal{L}^{\ psi,G}(M,p)$. The following is Theorem 1.7 from [Sp] and will be referred to as the holonomy theorem throughout the paper.
Theorem 2.1.
If $G$ is a connected and non-solvable Lie group, then there is a natural isomorphism $\mathcal{L}(M,p)\,\cong\,\mathcal{L}^{\ psi,G}(M,p)$ for every point $p$ in $M$.
In the case where the manifold $M$ is the euclidean space ${\mathbb{R}}^{n}$, the case considered in this paper, from now on we will simply denote by $\mathcal{L}_{p}$ the group $\mathcal{L}({\mathbb{R}}^{n},p)$.
@@ 2.2. Iterated Integrals and Signature
If $\gamma$ is a piecewise regular curve $\gamma$, Chen [Ch3] defines the formal power series
@@
in the noncommutative indeterminates $X_{1},\ldots,X_{n}$. This will be called the Chen signature map. Chen signature theorem can be stated as follows.
@@ 4. Holonomy and Signature
It is well known ([Ly], Theorem 2.1.1), ([Ner], Theorem 1.2) and [Re] that $\overline{G}_{n}$ is a Fréchet-Lie group.
Now, consider the trivial principal bundle $E\,=\,{\mathbb{R}}^{n}\times\overline{G}_{n}$ with connection $\nabla$ whose associated one-form is
$\omega_{\nabla}\,=-\,\sum_{i=1}^{n}\,X_{i}\,dx^{i},\qquad\omega_{\nabla}\in\Omega^{1}({\mathbb{R}}^{n},\,\overline{\mathfrak{g}}_{n}).$
The construction of the induced connection follows the same line as in the finite dimensional case, see for instance [Va].
@@
We prove that the holonomy $Hol_{\,\nabla,\,p}$ coincides with the Chen signature map (6) on loops.
By Theorem 2.1.2 in [Ly], the signature belongs to $\overline{G}_{n}$. We have to show that the signature is the holonomy of the connection form $\omega_{\nabla}$. Recall that $u$ is a horizontal curve in $E$ iff
(7) $u^{\prime}(t)=-\omega_{\nabla}(\gamma^{\prime}(t))u(t).$
By the argument in [Lo] (see also [Ner]), we obtain that the solution satisfying $u(0)=1$ is a series coinciding with the signature. Because of the Chen signature Theorem 2.2, the holonomy $Hol_{\,\nabla,\,p}$ is a monomorphism and the result follows. ∎
@@
Now notice that for every real $t$, we have $\omega_{\nabla}(\gamma^{\prime}(t))\in{\mathcal{T}}_{n}^{\xi}$. Therefore, the first part of Lemma 2.5 in [Ner] gives that the solution $u(t)$ of equation (7) belongs to the group $\mbox{Fr}_{n}^{\,\xi}$ and we have the result. ∎
@@
5. Geometric proof of Chen Signature Theorem

### v4

@@ Abstract.
We show that there is a topology on certain groups of loops in Euclidean space such that these groups are embedded in a Fréchet-Lie group which is the structural group of a principal bundle with connection whose holonomy coincides with the Chen signature map. We also give an alternative geometric new proof of the Chen signature theorem and a generalization of this theorem in classes strictly containing the one originally considered by Chen.
@@ 1. Introduction
In this paper we consider the relationship between holonomy maps of principal bundles with a connection and the signature map (see (4)) defined by Chen ([Ch1], [Ch2], [Ch3]).
If $\gamma$ is a piecewise regular curve in ${\mathbb{R}}^{n}$, Chen defines the formal power series
$\Theta(\gamma)\,=\,1\,+\,\sum^{\infty}_{p=1}\,\ \sum_{i_{1},\,i_{2},\,\ldots,\,i_{p}}\ \,\int_{\gamma}dx_{i_{1}}dx_{i_{2}}\ldots dx_{i_{p}}\ X_{i_{1}}X_{i_{2}}\ldots X_{i_{p}}$
in the noncommutative indeterminates $X_{1},\ldots,X_{n}$, where the coefficients are iterated integrals defined in section 2.2. This will be called the Chen signature map, and it is a homomorphism with respect to concatenation of curves. The Chen signature theorem (See Theorem 2.2 below) states that if $\Theta(\gamma)=1$ then $\gamma$ is a retraceable loop i.e. a trivial loop under the equivalence relation generated by the identifications $\alpha aa^{-1}\beta\sim\alpha\beta$. More recently (see [Ly, BGLY, HL, LX]) the signature map was generalized for curves with less regularity, and it was shown that the kernel of the signature map on paths with bounded variation consists of tree-like paths (see section 3).
@@
Holonomy maps appear in topology and differential geometry and they trace back to the well known theory of representations of the fundamental group and constructions of local systems over a manifold by the holonomy map of a principal bundle with a flat connection. A natural question is how to extend this construction to arbitrary connections. This question leads directly to the loop space of the manifold, an object that will be defined shortly  [^1] Do not confuse this space with the one defined in algebraic topology. The ambiguity in the names will disappear after we introduce the right notation.. In contrast to the fundamental group of a manifold, the loop space we are referring to here will not be a purely topological object. Specifically, while the fundamental group of a manifold coincides with that computed in the category of smooth curves, the loop space of the manifold, and therefore the results concerning this space, will depend on the category of curves where it is defined.
We show in Theorem 1.2 that the Chen signature map can be regarded as the holonomy map of a suitable connection on a principal bundle. The description of this very natural and simple geometric interpretation is one of the main aims of this note. The other one is to give a geometric proof of the Chen signature theorem, in Theorem 2.2, which is based on Theorem 1.7 of [Sp]. This approach also allows us to directly obtain a major ingredient in the original proof of the Chen signature theorem, namely Lemma 5.1 whose content is similar to the Fundamental Lemma 3.5 in [Ch3]. Below we expand on these ideas in more detail, including the precise statements of the main results and some extensions, as well as a discussion of the different loop spaces we consider.
@@
Consider a point $p$ in a manifold $M$. We say that a class of loops based at $p$ is concatenable if the concatenation of two loops in the class belongs to the class. Examples of classes with this property are the class of piecewise smooth loops $\Omega^{ps}(M,p)$, the class of piecewise smoothly immersed loops $\Omega^{psi}(M,p)$ or the class of piecewise analytic loops $\Omega^{pa}(M,p)$. By identifying loops of a concatenable class by reparameterization, we obtain associativity of the operation but in general $\alpha\,\alpha^{-1}\neq c$ where $c$ denotes the constant loop. A further equivalence relation is needed, and we consider the following three. Firstly, a natural relation is the one finitely generated by $\alpha aa^{-1}\beta\sim\alpha\beta$. This will be called the retrace relation. The corresponding groups are denoted by $\mathcal{L}^{\ ps,ret}(M,p)$, $\mathcal{L}^{\ psi,ret}(M,p)$ or $\mathcal{L}^{\ pa,ret}(M,p)$ depending on the choice of the class of curves. Then there is the relation by thin or zero area homotopies, that we define in Section 2. We denote the resulting groups by $\mathcal{L}^{\ ps,thin}(M,p)$, $\mathcal{L}^{\ psi,thin}(M,p)$ or $\mathcal{L}^{\ pa,thin}(M,p)$ respectively.
@@
For this work it will be important to consider the difference between the groups obtained by these equivalence relations. We quote the known facts about this in Section 2, but we remark one result here: In [Me], Meneses provides a non-retraceable thin loop in the class of piecewise smoothly loops, showing in particular that the group $\mathcal{L}^{\ ps,ret}(M,p)$ is a proper extension of $\mathcal{L}^{\ ps,thin}(M,p)$. We show that no such counterexample is possible in the class of piecewise smoothly immersed loops.
Actually, in section 3 we prove the following proposition.
@@
From now on, unless otherwise specified, we will denote the groups in the previous proposition simply by $\mathcal{L}(M,p)$ and by $\mathcal{L}_{p}$ in the case the manifold is the Euclidean space ${\mathbb{R}}^{n}$.
In our next result we construct a principal bundle whose holonomy map is the signature on loops. Let $\overline{G}_{n}$ be the Fréchet-Lie group corresponding to the completion of the real free Lie algebra generated by $X_{1},\ldots,X_{n}$. (See section 4 for more details and references).
@@
There is a $\overline{G}_{n}$-principal bundle with connection $(E,\,{\mathbb{R}}^{n},\,\overline{G}_{n},\,\pi,\,\theta)$ such that for any $p$ in ${\mathbb{R}}^{n}$, the holonomy map
$Hol_{\,\theta,\,p}:\,\mathcal{L}_{p}\,\hookrightarrow\,\overline{G}_{n}$
is a monomorphism and it coincides with the Chen signature map on $\mathcal{L}_{p}$.
For this work we consider principal bundles with left actions of the structure group. This is not the most common convention but it is what makes the holonomy a group homomorphism. When the action is on the right, the resulting holonomy reverses the order of the product. Chen [Ch2] also defines an alternative version of the signature $\Theta^{*}$, satisfying $\Theta^{*}(\alpha\beta)=\Theta^{*}(\beta)\Theta^{*}(\alpha)$. Taking the action on the right would lead to $\Theta^{*}$ as holonomy (see Remark 1). One feature of our approach is that it provides simple geometric characterizations of both signatures.
In [Ner], the author introduces for $\xi>0$ a family of Banach-Lie groups $\mbox{Fr}_{n}^{\,\xi}\subset\overline{G}_{n}$ allowing us to strengthen the above result:
@@
For every loop $\gamma$ in $\mathcal{L}_{p}$, the element $Hol_{\,\theta,\,p}(\gamma)\,=\,\Theta(\gamma)$ belongs to the Banach-Lie group $\mbox{Fr}_{n}^{\,\xi}$ for every $\xi>0$.
In section 5 we give a geometric alternative new proof of the Chen signature theorem, quoted as Theorem 2.2 here. Our proof is based on Theorem 1.7 of [Sp] which states that $\mathcal{L}^{\ psi,ret}(M,p)\ \cong\ \mathcal{L}^{\ psi,G}(M,p)$ for connected and non-solvable $G$. This approach allows us to obtain directly a result (Lemma 5.1) whose content is similar to the Fundamental Lemma 3.5 in [Ch3]. The proof of Theorem 2.2 also shows that the Chen signature theorem also holds for piecewise smooth curves, replacing the result in [Sp] by a result in [Tl] giving that $\mathcal{L}^{\ ps,thin}(M,p)\ \cong\ \mathcal{L}^{\ ps,G}(M,p)$ for $G$ semisimple. This stronger version of the signature theorem implies in turn the following theorem.
Theorem 1.4.
@@
As an immediate corollary of the previous theorem we have the following complement to Theorem 1.2.
@@ 2.1. Loops and Holonomy
Another equivalence relation in ${\mathcal{O}}^{psi}(M)$ is given by the finite composition of thin homotopies. A thin homotopy between two curves $\alpha$ and $\gamma$ is a homotopy $\eta:[0,1]^{2}\rightarrow M$ such that its image is contained in the union of the images of the curves, that is
@@
Denote the quotient under this relation by ${\mathcal{E}}^{\,psi,ret}(M)$ and let $\mathcal{L}^{\ psi,thin}(M,p)$ be the classes based at $p$. Note that, as before, it is a group under concatenation whose neutral element is the equivalence class of $e_{p}$, the constant path at $p$.
@@
In [Tl], Tlas proves a similar result in the concatenable class of $C^{1}$ loops based at $p$ with zero derivative at $p$. He proves that if the Lie group $G$ is semisimple then a pair of loops in this class are $G$-holonomy related if and only if they are related by a thin homotopy. In the class of piecewise smooth loops, Tlas result is equivalent to the following isomorphism under thin homotopy relation (see Remark $4$ in [Me])
Along the lines of this result, Meneses in [Me] comments that the analog result of (1) claimed in [Sp] is not valid in the class of piecewise smooth loops and he gives a counterexample of a thin and non-retraceable loop, hence
One of the difficulties with piecewise smooth loops, in contrast with piecewise analytic ones, is that they can intersect in very complicated ways. This led Baez and Sawin in [BS] to consider instead the class of piecewise smoothly immersed loops $\Omega^{psi}(M,p)$. In this class they develop the technology of tassels and webs. A careful inspection into the work in [Sp], which uses the technology developed in [BS], shows that the claimed result holds for the class of piecewise smoothly immersed loops.
The following is Theorem 1.7 from [Sp] and will be referred to as the holonomy theorem throughout the paper.  [^2] We do not know why Spallanzani did not write the piecewise immersion hypothesis.
Theorem 2.1.
If $G$ is a connected and non-solvable Lie group, then there is a natural isomorphism $\mathcal{L}^{\ psi,ret}(M,p)\,\cong\,\mathcal{L}^{\ psi,G}(M,p)$ for every point $p$ in $M$.
@@
In view of this isomorphism, we will simply denote these groups by $\mathcal{L}(M,p)$. In the case where the manifold $M$ is the Euclidean space ${\mathbb{R}}^{n}$, the case considered in this paper, from now on we will simply denote by $\mathcal{L}_{p}$ the group $\mathcal{L}({\mathbb{R}}^{n},p)$.
@@ 2.2. Iterated Integrals and Signature
Recall from the introduction that if $\gamma$ is a piecewise regular curve $\gamma$, Chen [Ch3] defines the formal power series
@@
in the noncommutative indeterminates $X_{1},\ldots,X_{n}$. This will be called the Chen signature map. Chen proved that this map is a faithful representation of the group of loops $\mathcal{L}^{\ psi,ret}({\mathbb{R}}^{n},p)$ [Ch1, Ch2, Ch3]. For other results concerning iterated integrals, see [Ch4, Ch5]. In recent years this theory had a resurgence in the context of stochastic processes, bounded variation paths, and rough paths [Ly, BGLY, HL, LX].
The Chen signature theorem can be stated as follows.
@@ 4. Holonomy and Signature
It is well known ([Ly], Theorem 2.1.1), ([Ner], Theorem 1.2) and [Re] that $\overline{G}_{n}$ is a Fréchet-Lie group. Consider the trivial left principal bundle $E\,=\,{\mathbb{R}}^{n}\times\overline{G}_{n}$ and $\omega\in\Omega^{1}({\mathbb{R}}^{n},\,\overline{\mathfrak{g}}_{n})$ given by
$\omega=-\,\sum_{i=1}^{n}\,X_{i}\,dx^{i}.$
This form induces a connection one-form $\theta\in\Omega^{1}(E,\,\overline{\mathfrak{g}}_{n})$ given by
$\theta_{(p,g)}(v,X)=g\omega_{p}(v)g^{-1}+Xg^{-1}$
where $p,v\in{\mathbb{R}}^{n}$, $g\in\overline{G}_{n}$ and $X\in\overline{\mathfrak{g}}_{n}$. It is straightforward to check that this form is left invariant.
Remark 1.
The action of $\overline{G}_{n}$ on $E$ is on the left. Chen [Ch2] also defines an alternative version of the signature $\Theta^{*}$, satisfying $\Theta^{*}(\alpha\beta)=\Theta^{*}(\beta)\Theta^{*}(\alpha)$. Taking the action on the right would lead to $\Theta^{*}$ as holonomy.
@@
We prove that the holonomy $Hol_{\,\theta,\,p}$ coincides with the Chen signature map (4) on loops.
By Theorem 2.1.2 in [Ly], the signature belongs to $\overline{G}_{n}$. We have to show that the signature is the holonomy of the connection form $\theta$. Recall that $(\gamma,u)$ is a horizontal curve in $E$ iff for each $t$ we have $\theta(\gamma^{\prime}(t),u^{\prime}(t))=0$, i.e.
(5) $u^{\prime}(t)=-u(t)\omega(\gamma^{\prime}(t)).$
If $u$ is the solution satisfying $u(0)=1$ we obtain
(6) $u(t)=1+\sum_{i_{1}}\int_{t_{1}<t}u(t_{1})dx_{i_{1}}(\gamma^{\prime}(t_{1}))dt_{1}\;X_{i_{1}}.$
We write $u(t_{1})$ according to (6) integrating in $t_{2}$ which gives
(7) $\begin{split}u(t)&=1+\sum_{i_{1}}\int_{t_{1}<t}dx_{i_{1}}(\gamma^{\prime}(t_{1}))dt_{1}\;X_{i_{1}}+\\ &+\sum_{i_{2}i_{1}}\int_{t_{2}<t_{1}<t}u(t_{2})dx_{i_{2}}(\gamma^{\prime}(t_{2}))dx_{i_{1}}(\gamma^{\prime}(t_{1}))dt_{2}dt_{1}\;X_{i_{2}}X_{i_{1}}.\end{split}$
Repeating this procedure (see also [Lo], [Ner]) we obtain the series
coinciding with the signature. Because of the Chen signature Theorem 2.2, the holonomy $Hol_{\,\theta,\,p}$ is a monomorphism and the result follows. ∎
This procedure also gives that the signature is a group element.
@@
Now notice that for every real $t$, we have $\omega(\gamma^{\prime}(t))\in{\mathcal{T}}_{n}^{\xi}$. Therefore, the first part of Lemma 2.5 in [Ner] gives that the solution $u(t)$ of equation (5) belongs to the group $\mbox{Fr}_{n}^{\,\xi}$ and we have the result. ∎
@@
5. Geometric proof of the Chen Signature Theorem